OpenStudy (anonymous):

I need help setting up this problem: A flat rectangular piece of aluminum has a perimeter of 60 inches. The length is 8 inches longer than the width. Find the width.

OpenStudy (jtvatsim):

OK, so the perimeter is P = 60, but you also know that the perimeter of something is just P = 2L + 2W. You also know that L = W + 8. So plug in the values to get: 60 = 2(W + 8) + 2W and solve for W.

OpenStudy (anonymous):

sweet thanks

OpenStudy (anonymous):

perimeter of rectangle is P=2l+2W not something :).

OpenStudy (jtvatsim):

@sami-21 of course! :)